# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 5.1 Invariance inside chambers
Similarly to the proof of Theorem 3.2, we start with the verification of the constancy of $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ along intervals $t^{\prime}<t<t^{\prime\prime}$ such that $\{\widehat{\boldsymbol{w}}(t)\}_{t^{\prime}<t<t^{\prime\prime}}$ lies in the same chamber of $\operatorname{Men}_{2}(\Delta)$. To this end, it is enough to show that the projection ${\mathcal{M}}_{1,n}^{\;{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}\to(t^{\prime},t^{\prime\prime})$ has no critical points (cf. the proof of Theorem 3.2). This requirement amounts to the following statement (cf., Lemma 4.1 and [23, Lemma 2.3]).
@@
Let $\xi=[{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{1,n}^{\;{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$, where ${\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$. Then
$H^{0}\left({\mathbb{E}},{\mathcal{N}}_{\Phi}\left(-p_{0}-\sum_{\sigma\neq\sigma_{0},i\neq i_{0}}2k^{\sigma}_{i}p^{\sigma}_{i}-(2k^{\sigma_{0}}_{i_{0}}-1)p^{\sigma_{0}}_{i_{0}}\right)\right)=0.$ (22)
Proof. We proceed along the lines of the proof of [23, Lemma 2.3]. Assume that $H^{0}\neq 0$ in (22). Then, there exists a real curve $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C\}$, where $C={\Phi}_{*}({\mathbb{E}})$, which intersects $C$
@@ 5.2 Walls in the space of constraints
The set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\times{\mathcal{Q}}$ satisfying
@@
is a dense semialgebraic subset of full dimension $\dim\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\times{\mathcal{Q}}=n+1$. The complement is the union of finitely many semialgebraic strata of codimension $\geq 1$. Since the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ is generic, it avoids strata of $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\times{\mathcal{Q}}$ of codimension $\geq 2$ and intersects strata of codimension one only in their generic points.
Now, we study the strata $S$ of the dimension $n$ in $(\operatorname{Men}_{\mathbb{R}}^{\tau}\times{\mathcal{Q}})\setminus\operatorname{Men}_{1}(\Delta)$ in terms of the geometry of elements $\xi\in\overline{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))\setminus{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))$, $0\leq t\leq 1$.
@@
    $\widehat{C}$ is a reducible curve of arithmetic genus $1$ with a component mapped onto a toric divisor;
@@
    $\widehat{C}$ is a connected curve of arithmetic genus $1$ either with at least three irreducible components, or with two rational irreducible components and no other components.
(2) If $\xi=[{\Phi}:(\widehat{C},\widehat{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$, where $\widehat{\boldsymbol{w}}$ is a generic element in an $n$-dimensional stratum in $(\operatorname{Men}_{\mathbb{R}}^{\tau}\times{\mathcal{Q}})\setminus\operatorname{Men}_{1}(\Delta)$, then $\xi$ is of one of the following types:
(2a) either $w_{0}(t^{*})$ is a singular point of the curve $C={\Phi}_{*}(\widehat{C})$, and the following holds:
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, the curve $\widehat{C}$ is smooth elliptic, ${\Phi}$ is an immersion onto a curve $C$ that is smooth along the toric divisors, while $w_{0}(t^{*})$ is a center of at least one real local branch;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, the curve $\widehat{C}$ is smooth elliptic, ${\Phi}$ is birational onto its image, but not an immersion; furthermore, $C={\Phi}(\widehat{C})$ is smooth at $w_{0}(t^{*})$ and at each point $w_{i}^{\sigma}(t^{*})$ with $k_{i}^{\sigma}\geq 2$, and is unibranch at each point $w_{i}^{\sigma}(t^{*})$ with $k_{i}^{\sigma}=1$;
@@
    $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$ for some $\sigma\in P^{1}_{\Delta}$ and $i\neq j$, the curve $\widehat{C}$ is smooth elliptic, and the map ${\Phi}$ is an immersion such that the point $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$ is a center of one or two smooth branches;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{2}$ is a smooth elliptic curve, the intersection $\widehat{C}_{1}\cap\widehat{C}_{2}$ consisting of one point $p$; the map ${\Phi}:\widehat{C}_{1}\to{\operatorname{Tor}}(P_{\Delta})$, $s=1,2$, is either an immersion, smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$, or a multiple covering of a line intersecting only two toric divisors, while these divisors correspond to opposite parallel sides of $P_{\Delta}$ and the intersection points with these divisors are ramification points of the covering; the map ${\Phi}:\widehat{C}_{2}\to{\operatorname{Tor}}(P_{\Delta})$ is an immersion, smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$; the point $p$ either is mapped to ${\operatorname{Tor}}(P_{\Delta})^{\times}$, and then the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ intersect only in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ and each of their intersection point is an ordinary node, or $p$ is mapped to some point $w_{i}^{\sigma}(t^{*})$, in which case the curves $C_{i}={\Phi}(\widehat{C}_{i})$, $i=1,2$, intersect ${\operatorname{Tor}}(\sigma)$ with even multiplicities, and they do not have other common point in $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}$ is an irreducible rational curve with one node $p$, the map ${\Phi}$ is an immersion that sends $p$ to some point $w_{i}^{\sigma}(t^{*})$, center of two smooth branches intersecting ${\operatorname{Tor}}(\sigma)$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, where $l_{1}+l_{2}+1=k_{i}^{\sigma}$, and intersecting each other with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$, furthermore, each of the other points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is a center of one smooth branch;
@@
    $\widehat{C}$ is a smooth elliptic curve, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n=4$ points, the curve $C={\Phi}({\mathbb{P}}^{1})$ is rational and smooth at $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and ${\Phi}:\widehat{C}\to C$ is a double covering ramified at $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$.
Proof. (1i) In such a case, we must have $\widehat{C}=\widehat{C}^{\prime}\cup\widehat{C}^{\prime\prime}$, where $\widehat{C}^{\prime}$ is a connected curve of arithmetic genus one suhc that $\widehat{C}^{\prime}$ is mapped onto the union of all toric divisors, and $\widehat{C}^{\prime\prime}$ is a non-empty (due to $w_{0}\in{\operatorname{Tor}}(P_{\Delta})^{\times}$) union of connected curves of arithmetic genus zero, each one joined with $\widehat{C}^{\prime}$ in one point (cf. the proof of Lemma 4.2(1i)). Let $\widehat{C}^{\prime\prime}_{0}$ be a connected component of $\widehat{C}^{\prime\prime}$ such that $w_{0}(t^{*})\in{\Phi}(\widehat{C}^{\prime\prime}_{0})$. Note that all local branches of ${\Phi}:\widehat{C}^{\prime\prime}_{0}\to{\operatorname{Tor}}(P_{\Delta})$ centered on toric divisors, except for ${\Phi}:(\widehat{C}^{\prime\prime}_{0},p)\to{\operatorname{Tor}}(P_{\Delta})$, where $p=\widehat{C}^{\prime}\cap\widehat{C}^{\prime\prime}_{0}$, are, in fact, centered at some points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. Observe that either $C^{\prime\prime}_{0}$ intersects at least three toric divisors of ${\operatorname{Tor}}(P_{\Delta})$, or it intersects two toric divisors corresponding to parallel sides of $P_{\Delta}$. However, in the both cases we reach a contradiction with (AQC).
@@
(2a) Let $w_{0}(t^{*})$ be a singular point of $C$. Then, for the dimension reason, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ must be a generic element of $\operatorname{Men}_{\mathbb{R}}^{\tau}$. This implies, in particular, that $\widehat{C}$ is irreducible and $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points. Furthermore, $\widehat{C}$ cannot be a rational curve with a node. Thus, $\widehat{C}$ is a smooth elliptic curve. Then (see Lemma 2.4) we derive that ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is an immersion and $C$ is smooth along the toric divisors. Since $w_{0}(t^{*})$ turns into a real smooth point in the deformation along the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$, it must be a center of at least one real local branch.
@@
Suppose that $\widehat{C}$ is a smooth elliptic curve and $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points. Then, we obtain the claim (2ii) due to Lemma 2.4 and the claim (1ii) above.
@@
Suppose that some of the points of the sequence $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ coincide. For the dimension reason, we immediately get that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, and all these points must be in general position subject to the unique Menelaus relation (1). Hence, $\widehat{C}$ must be irreducible. Moreover, $\widehat{C}$ cannot be rational, since otherwise, by Lemma 2.4, one would get the total dimension of the considered stratum $\leq(n-2)+1=n-1<n$, a contradiction. Thus, $\widehat{C}$ is a smooth elliptic curve. We have $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=w$ for some $\sigma\subset\partial P$ and $i\neq j$. If $C$ is unibranch at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, then, by Lemma 2.4, the curve $C$ is immersed and smooth along the toric divisors. Otherwise, the curve $C$ has two local branches at $w$. Moreover, it must be unibranch at each point $w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\neq w$ and, in addition, smooth if $k_{i^{\prime}}^{\sigma^{\prime}}\geq 2$ due to Lemma 2.4 and the claim (1ii). Let us show that $C$ is immersed. Fixing the position of $w_{0}(t^{*})$ and the position of one more point $w^{\prime}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$, we obtain a family of dimension $\geq n-2\geq 2$; hence, [13, Inequality (5) in Lemma 2.1] applies:
$c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})\geq(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-n+2)+1+\sum_{B}({\operatorname{ord}}B-1)+(n-3)$
$=c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+\sum({\operatorname{ord}}B-1)\ ,$
where $B$ runs over all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$. That is, $C$ is immersed, and we finally fit the requirements of claim (2iii).

### v2

@@ 5.1 Invariance inside chambers
Similarly to the proof of Theorem 3.2, we start with the verification of the constancy of $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ along intervals $t^{\prime}<t<t^{\prime\prime}$ such that $\{\widehat{\boldsymbol{w}}(t)\}_{t^{\prime}<t<t^{\prime\prime}}$ lies in the same chamber of $\operatorname{Men}_{1}(\Delta)$. To this end, it is enough to show that the projection ${\mathcal{M}}_{1,n}^{\;{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}\to(t^{\prime},t^{\prime\prime})$ has no critical points (cf. the proof of Theorem 3.2). This requirement amounts to the following statement (cf., Lemma 4.1 and [29, Lemma 2.3]).
@@
Let $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{1,n}^{\;{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$, where ${\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$. Then,
$H^{0}\left(\widehat{C},{\mathcal{N}}_{\Phi}\left(-p_{0}-\sum_{\sigma\neq\sigma_{0},i\neq i_{0}}2k^{\sigma}_{i}p^{\sigma}_{i}-(2k^{\sigma_{0}}_{i_{0}}-1)p^{\sigma_{0}}_{i_{0}}\right)\right)=0.$ (30)
Proof. We proceed along the lines of the proof of [29, Lemma 2.3]. Assume that $H^{0}\neq 0$ in (30). Then, there exists a real curve $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C\}$, where $C={\Phi}_{*}(\widehat{C})$, which intersects $C$
@@ 5.2 Walls in the space of constraints
The set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)\times{\mathcal{Q}}$ satisfying
@@
is a dense semialgebraic subset of full dimension $\dim\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)\times{\mathcal{Q}}=n+1$. The complement is the union of finitely many semialgebraic strata of codimension $\geq 1$. Since the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ is generic, it avoids strata of $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)\times{\mathcal{Q}}$ of codimension $\geq 2$ and intersects strata of codimension one only in their generic points. Note that the image ${\operatorname{pr}}(S)\subset|{\mathcal{L}}_{P_{\Delta})}|$ of ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$ in the linear system $|{\mathcal{L}}_{P_{\Delta}}|$ when $\widehat{\boldsymbol{w}}$ ranges over a stratum $S$ of codimension one, has dimension $n-1$.
Now, we study the strata $S$ of the dimension $n$ in $(\operatorname{Men}_{\mathbb{R}}^{\rho}\times{\mathcal{Q}})\setminus\operatorname{Men}_{1}(\Delta)$ in terms of the geometry of elements $\xi\in\overline{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))\setminus{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))$, $0\leq t\leq 1$. In view of the results of Section 5.1, we can ignore the strata along which ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$. The following lemma classifies other strata.
@@
    $\widehat{C}$ is a reducible curve of arithmetic genus $1$ having a component mapped onto a toric divisor;
@@
    $\widehat{C}$ is a connected curve of arithmetic genus $1$ either having at least three irreducible components, or with two rational irreducible components and no other components.
(2) If $\xi=[{\Phi}:(\widehat{C},\widehat{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overline{\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ when $\widehat{\boldsymbol{w}}(t^{*})$ does not satisfy the conclusions of Lemma 2.4, then $\xi$ is of one of the following types:
(2a) either $w_{0}(t^{*})$ is a singular point of the curve $C={\Phi}(\widehat{C})$, and the following holds:
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, the curve $\widehat{C}$ is smooth elliptic, ${\Phi}$ is an immersion onto a curve $C$ which is smooth along the toric divisors, while $w_{0}(t^{*})$ is a center of at least one real local branch;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}$ is a smooth elliptic curve, the map ${\Phi}$ is birational onto its image $C$ but not an immersion: it may have singular local branches centered in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ as well as singularities of type $A_{2m}$ at some points $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, where $m\geq k_{i}^{\sigma}$;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, in particular, $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=:w$ for some $\sigma\in P^{1}_{\Delta}$ and $i\neq j$, the curve $\widehat{C}$ is smooth elliptic, and the map ${\Phi}$ is birational onto its image $C$; the curve $C$ is immersed outside $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$ and may have singularities of type $A_{2m}$ at some points $w_{i^{\prime}}^{\sigma^{\prime}}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$, while at $w$, the curve $C$ either is smooth or has two smooth branches;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{2}$ is a smooth elliptic curve, the intersection $\widehat{C}_{1}\cap\widehat{C}_{2}$ consists of one point $p$; the map ${\Phi}:\widehat{C}_{1}\to{\operatorname{Tor}}(P_{\Delta})$, $s=1,2$, is either an immersion, smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$, or a multiple covering of a line intersecting only two toric divisors, while these divisors correspond to opposite parallel sides of $P_{\Delta}$ and the intersection points with these divisors are ramification points of the covering; the map ${\Phi}:\widehat{C}_{2}\to{\operatorname{Tor}}(P_{\Delta})$ is an immersion, smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$; the point $p$ either is mapped to ${\operatorname{Tor}}(P_{\Delta})^{\times}$, and then the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ intersect each other only in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ and each of their intersection points is an ordinary node, or $p$ is mapped to some point $w_{i}^{\sigma}(t^{*})$, in which case the curves $C_{i}={\Phi}(\widehat{C}_{i})$, $i=1,2$, intersect ${\operatorname{Tor}}(\sigma)$ with even multiplicities, and they do not have other common point in $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$;
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}$ is an irreducible rational curve with one node $p$, the map ${\Phi}$ is an immersion sending $p$ to some point $w_{i}^{\sigma}(t^{*})$, center of two smooth branches intersecting with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$ and intersecting ${\operatorname{Tor}}(\sigma)$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, where $l_{1}+l_{2}+1=k_{i}^{\sigma}$; furthermore, each of the other points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is a center of one smooth branch;
@@
    $\widehat{C}$ is a smooth elliptic curve, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n=4$ points, the curve $C={\Phi}({\mathbb{P}}^{1})$ is rational and smooth at $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and ${\Phi}:\widehat{C}\to C$ factors through the double covering $\widehat{C}\to C^{\nu}\simeq{\mathbb{P}}^{1}$ ramified at four points projecting to $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, where $C^{\nu}\to C$ is the normalization.
Proof. (1i) In this case, we must have $\widehat{C}=\widehat{C}^{\prime}\cup\widehat{C}^{\prime\prime}$, where $\widehat{C}^{\prime}$ is a connected curve of arithmetic genus one such that $\widehat{C}^{\prime}$ is mapped onto the union of all toric divisors, and $\widehat{C}^{\prime\prime}$ is a non-empty (due to $w_{0}\in{\operatorname{Tor}}(P_{\Delta})^{\times}$) union of connected curves of arithmetic genus zero, each one joined with $\widehat{C}^{\prime}$ in one point (cf. the proof of Lemma 4.2(1i)). Let $\widehat{C}^{\prime\prime}_{0}$ be a connected component of $\widehat{C}^{\prime\prime}$ such that $w_{0}(t^{*})\in{\Phi}(\widehat{C}^{\prime\prime}_{0})$. Note that all local branches of ${\Phi}:\widehat{C}^{\prime\prime}_{0}\to{\operatorname{Tor}}(P_{\Delta})$ centered on toric divisors, except for ${\Phi}:(\widehat{C}^{\prime\prime}_{0},p)\to{\operatorname{Tor}}(P_{\Delta})$, where $p=\widehat{C}^{\prime}\cap\widehat{C}^{\prime\prime}_{0}$, are, in fact, centered at some points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. Observe that either $C^{\prime\prime}_{0}$ intersects at least three toric divisors of ${\operatorname{Tor}}(P_{\Delta})$, or it intersects two toric divisors corresponding to parallel sides of $P_{\Delta}$. However, in the both cases we reach a contradiction with (AQC).
@@
(2a) Let $w_{0}(t^{*})$ be a singular point of $C$. Then, for the dimension reason, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ must be a generic element of $\operatorname{Men}_{\mathbb{R}}^{\rho}$. This implies, in particular, that $\widehat{C}$ is irreducible and $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points. Furthermore, $\widehat{C}$ cannot be a rational curve with a node. Thus, $\widehat{C}$ is a smooth elliptic curve. Then (see Lemma 2.4), we derive that ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is an immersion and $C$ is smooth along the toric divisors. Since $w_{0}(t^{*})$ turns into a real smooth point in the deformation along the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$, it must be a center of at least one real local branch.
@@
(2ii) Suppose that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}$ is a smooth elliptic curve, ${\Phi}:\widehat{C}\to C={\Phi}(\widehat{C})$ is birational, but not an immersion. To verify the conditions of item (2ii) of the lemma, it is enough to consider the case of $C$ having a singular branch at some point $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. Fixing the position of the point $w_{i}^{\sigma}$, we obtain a subfamily ${\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\subset{\operatorname{pr}}(S)$ of dimension $\geq n-2\geq 2$ and then apply [16, Inequality (5) in Lemma 2.1] (cf. (13)):
$-CK_{{\operatorname{Tor}}(P_{\Delta})}\geq\left(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+1\right)+(n-3)+({\operatorname{mult}}Q-1)$
$=-CK_{{\operatorname{Tor}}(P_{\Delta})}-2+({\operatorname{mult}}Q-1)\quad\Longrightarrow\quad{\operatorname{mult}}Q\leq 3\ ,$ (31)
where $Q$ is the local branch of $C$ at $w_{i}^{\sigma}$. It remains to show that ${\operatorname{mult}}Q=2$. Arguing by contradiction, assume that ${\operatorname{mult}}Q=3$, i.e., $Q$ is topologically equivalent to a singularity $y^{3}+x^{m}$, where $\gcd(3,m)=1$ and $m\geq 2k_{i}^{\sigma}\geq 4$. Without loss of generality, we can assume that the germ of ${\operatorname{pr}}(S)\subset|{\mathcal{L}}_{P_{\Delta}}|$ at $C$ is an equisingular family. According to items (a) and (b) in Section 2.3, we obtain the following upper bound completely analogous to (5):
$\dim{\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\leq h^{0}(C,{\mathcal{J}}_{Z/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}}=h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right),$ (32)
where
$\deg{\boldsymbol{d}}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+1)+2$
$\Longrightarrow\quad\deg({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=C^{2}-\deg{\boldsymbol{d}}=n-3\geq 1.$
By the Riemann-Roch formula, this yields
$\dim{\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\leq n-3$
contradicting the lower bound $\dim{\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\geq n-2$ pointed above.
@@
(2iii) Suppose that some of the points of the sequence $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ coincide. For the dimension reason, we immediately get that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, and all these points must be in general position subject to the unique Menelaus relation (1). Hence, $\widehat{C}$ must be irreducible. Moreover, $\widehat{C}$ cannot be rational, since otherwise, by Lemma 2.4, one would get the total dimension of the considered stratum $\leq(n-2)+1=n-1<n$, a contradiction. Thus, $\widehat{C}$ is a smooth elliptic curve. We have $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=w$ for some $\sigma\subset\partial P$ and $i\neq j$. If $C$ is unibranch at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, then, by Lemma 2.4, the curve $C$ is immersed and smooth along the toric divisors. Otherwise, the curve $C$ has two local branches at $w$. Moreover, it must be unibranch at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$.
Let us analyze possible singularities of $C$. Fixing the position of $w$ and some other point $w^{\prime}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$, we obtain a family ${\operatorname{pr}}(S^{fix}(w,w^{\prime}))\subset{\operatorname{pr}}(S)$ of dimension $\geq n-3\geq 1$; hence, [16, Inequality (5) in Lemma 2.1] applies and similarly to (13) yields
$-CK_{{\operatorname{Tor}}(P_{\Delta})}\geq\left(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+3\right)+(n-4)+\sum_{B}({\operatorname{mult}}Q-1)$
$=-CK_{{\operatorname{Tor}}(P_{\Delta})}-1+\sum_{B}({\operatorname{mult}}Q-1)\quad\Longrightarrow\quad\sum_{B}({\operatorname{mult}}Q-1)\leq 1\ ,$ (33)
where $B$ is the set of all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$. That means a singular branch of $C$ (if any) must be of order $2$. We show that $C$ is immersed outside $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$, and hence fits the claim of item (2iii).
Let $C$ have a singular local branch $Q$ centered in ${\operatorname{Tor}}(P_{\Delta})^{\times}$. Note that $\dim{\operatorname{pr}}(S^{fix}(w))\geq n-2$. Similarly to (32) we get
$\dim{\operatorname{pr}}(S^{fix}(w))\leq h^{0}(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}),$ (34)
where
$\deg{\boldsymbol{d}}_{1}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+2)+({\operatorname{mult}}Q-1)$
$\Longrightarrow\quad\deg(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=C^{2}-\deg{\boldsymbol{d}}_{1}=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})})$
$-(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+2)-({\operatorname{mult}}Q-1)=n-3\geq 1,$ (35)
which yields $\dim{\operatorname{pr}}(S^{fix}(w))\leq n-3$, and hence a contradiction with the lower bound $\dim{\operatorname{pr}}(S^{fix}(w))\geq n-2$.
Let $C$ have a singular local branch $Q$ centered at $w$. Then $\dim{\operatorname{pr}}(S^{fix}(w))\geq n-2$, whereas the computation literally coinciding with that in the preceding paragraph yields $\dim{\operatorname{pr}}(S^{fix}(w))\leq n-3$.